# Akshay Krishnan1, 2  Xinchen Yan1  Vincent Casser3  Abhijit Kundu1 1Google DeepMind  2Georgia Insitute of Technology  3Waymo

### v1

@@ 3.3 Color-conditioned depth and normal prediction
Figure 4: Text-conditioned generation: Color, depth and normal images generated jointly by our model from a single text prompt. The input text prompt is provided on the top. The generated color, depth, and normal are presented from top to bottom. Results show that Orchid can produce consistent appearance and geometry for a wide range of text prompts such as close-up object shots or views from complex layouts of indoor and outdoor environment.
The formulation in Section 3.2 allows us to generate color, depth, and normals directly from the joint latent space through $f_{\theta_{\text{LDM}}}(\mathbf{z}_{t},t)$. To further use Orchid to generate depth and surface normals jointly from an input image condition, we append an input image latent $\mathbf{z}^{\mathbf{x}}$ from the pretrained frozen color-only VAE as a condition signal for diffusion training through $f_{\theta_{\text{LDM}}}(\mathbf{z}_{t},t;\mathbf{z}^{\mathbf{x}})$. While this is in line with previous work on using diffusion priors for dense prediction [25, 18], our joint latent space allows direct generation of both depth and surface normals from image conditions. When used as an image-conditioned geometry estimator at inference, we drop text conditions (Figure 3b). The sampling procedure is detailed in Alg. 2.
Algorithm 2 Color-conditioned depth sampling
$\mathbf{z}^{\mathbf{x}}$ (color-space latent condition), $T$ (total denoising steps)
@@ 4.1 Implementation details
We use a convolutional encoder and decoder with a latent channel of 8 for our joint VAE, initialized from a VAE pretrained on datasets of color images alone. For the latent diffusion model, we use a transformer UNet similar to Stable Diffusion [41] pretrained on a large image-text dataset. The diffusion model uses v-prediction. We train both models on 512x512 image resolution. When training our text and color-conditioned diffusion models, we use a DDPM noise schedule [20] with 1000 steps. For diffusion sampling, we adopt a DDIM schudule with 100 steps for text-conditioned generation and 50 steps for image-conditioned generation. We use the zero terminal SNR schedule from [33].
Dataset: We combine many real and synthetic datasets with ground truth depth and surface normals to train our VAE: Omnidata [9], Virtual KITTI [14], Hypersim [40], and DIODE [53]. Additionally, we create a large-scale dataset with text, color, depth and normals by using large pretrained models to predict depth [60] and surface normal [22] from a text-image dataset. We use approximately 2.5M images with ground truth depth and surface normal, and about 100M images with pseudo ground truth.
@@ 4.2 Text-conditioned generation
Orchid is a first-of-its-kind model that can jointly generate color images along with their depth and normal from text prompts. We provide qualitative results of color, depth and normal generated by Orchid from a diverse set of text prompts in Figure 4. Assuming approximate intrinsics, we can use the predicted color, depth and surface normal to optimize 3D Gaussians using Gaussian Splatting [26]. We show an example of this in Figure 1. Additional results are provided in Appendix C.
@@
Table 2: Zero-shot monocular depth estimation: Comparison of our zero shot affine-invariant depth estimation accuracy to that of other methods. Our accuracy is comparable to that of [60] while being better than the other joint depth-normal method [13]. The first, second and third ranking methods are highlighted.
@@
Table 3: Zero-shot monocular surface normals estimation: Comparison of our color-conditioned normal estimation accuracy to state-of-the-art. Orchid is significantly better than other methods while jointly predicting both depth and normals. The first, second and third ranking methods are highlighted.
Figure 6: Monocular depth prediction: Orchid obtains more accurate depth compared to Marigold [25], as highlighted in the results shown above. Our model achieves better performance on far-range objects (row 1,2,4), small objects (row 5), depth-ambiguous scenes (row 6), and transparent occlusions (row 3).
Figure 7: Monocular surface normal prediction: A comparison of our jointly generated normal to that of color-conditioned prediction using Marigold [25]. Orchid produces more robust results on curved and deformable surfaces.
Dataset Marigold [25] GeoWizard [13] Orchid
NYUv2 0.102 0.122 0.040
KITTI 0.146 0.324 0.082
Table 4: Depth-normal inconsistency: We show the mean error $e_{\text{depth\_normal}}$ ($\downarrow$) in the table. Orchid’s color-conditioned joint depth and normal estimates are significantly more consistent than those produced by GeoWizard or separate Marigold models.
@@
Figure 9: Joint color-depth-normal inpainting: Given a masked region in color-depth-normal data (left column shows masked color), our model inpaints them jointly using an approach like [35].
@@ 4.3 Monocular depth and normal prediction
The joint color-depth-normal prior learned by Orchid can also be leveraged for monocular depth and normal prediction with color-conditioned fine-tuning (Section 3.3). We evaluate the accuracy of the depth and normal jointly estimated by Orchid from color images. We follow the protocol in recent approaches and evaluate the zero-shot performance on standard benchmarks we do not train on. In addition, we present qualitative results in Figure 5. As illustrated in the figure, our model can produce high-fidelity and consistent depth and normal estimation from in-the-wild color images that contain single objects, indoor, and outdoor complex scene environment.
@@
We evaluate the accuracy of the affine-invariant depth estimated by Orchid on four held-out datasets: NYUDepthv2 [36], KITTI [15], ETH3D [50] and ScanNet [7]. We compare Orchid’s performance to other zero-shot depth estimation baselines. This includes feed-forward momdels [39, 22, 60] as well as models that learn a conditional depth estimation prior by finetuning image generative models [25, 13]. Note that all our baselines (except [13]) are explicitly trained for color-conditioned depth prediction while Orchid jointly predicts depth and normals. As demonstrated in Table 2, Orchid’s depth estimates are comparable to the state-of-the-art method [60], while being better than other diffusion-based baselines. As illustrated in Figure 6, our method produces more accurate depth than Marigold [25] on far-range objects, small objects, and depth-ambiguous scenes. Notably, Orchid achieves better monocular depth estimation performance than Geowizard [13], the only baseline that jointly predicts depth and surface normals.
@@
We also evaluate Orchid’s normal estimation accuracy on four held-out datasets: NYUv2 [36], ScanNet [7], iBims-1 [28] and Sintel [4]. As shown in Table 3, Orchid is significantly better than baselines at estimating surface normal. This highlights the significance of training a VAE that explicitly encodes color, depth and normal into a joint latent on a large scale dataset, as opposed to other diffusion baselines [25, 13, 18] that use a frozen color-space VAE to encode and decode depth and normal. We visualize the comparisons to Marigold-normal [25] in Figure 7, where our method has stronger normal prediction performances on curved and deformable surfaces.
@@
We evaluate the consistency of the depth and surface normal produced by Orchid and compare it to other diffusion baselines that either use separate model weights or switch embedding to predict depth and normal [25, 13]. We align the affine-invariant depth $\mathbf{\hat{d}}$ with ground truth to get metric depth predictions $\mathbf{d}$, which we use with camera intrinsics to compute a 3D pointcloud $\mathbf{p}$. We use the pointcloud to estimate normals $\mathbf{\hat{n}}=\nabla_{x}\mathbf{p}\times\nabla_{y}\mathbf{p}$ from depth, and compute its inconsistency $e_{\text{depth\_normal}}=(1-\mathbf{\hat{n}}\cdot\mathbf{n})/2$ with the estimated normals $\mathbf{n}$. As shown in Table 4, we find that Orchid is able to leverage the joint latent space to predict depth and normal that are significantly more consistent. We visualize the depth-normal inconsistency error map in Figure 8. These results clearly demonstrate the benefit of our joint generative model of appearance and geometry in downstream applications.
@@ 4.4 Joint color-depth-normal inpainting
The joint color-depth-normal generative prior learned by Orchid can be used for unconditional tasks that combine appearance and geometry, such as jointly inpainting both color and geometry. Given color, depth and normals for a frame, we mask all of them for the region to be inpainted (input in Figure 9). Following the approach from RePaint [35], we use Orchid as an unconditional inpainting prior to generate inpainted regions, as shown in Figure 9. We find that our model is able to generate plausible inpainted regions with consistent color, depth and normals, without any additional training. This is a unique capability that results from Orchid’s joint prior. Please refer to Appendix C.4 for more inpainting results.
@@ 4.5 Ablations
We ablated the impact of some choices we make in training Orchid’s VAE and latent diffusion model. Please find additional ablations in Appendix B.
Distillation loss: We use a distillation loss to encourage our VAE’s latents to be close in distribution to the original LDM’s latents. We find that dropping the distillation loss during VAE training makes the diffusion training forget some priors learned from image pretraining, generating qualitatively worse images. We observe a reduced CLIP image-text similarity by $0.7$ and increased FID by $1.7$.
Use of pseudo labels: Unlike previous methods that rely solely on real or synthetic training data, we use a large text-image dataset with depth and normals from pretrained models. We find that this improves the VAE’s depth predictions by 1.3 AbsRel and normals by $7.7\%$ ($11.25\degree$ error).
@@ 5 Conclusion
In this work, we presented Orchid, a joint appearance and geometry diffusion prior that encodes color, depth, and surface normal in an unified latent space. This allows Orchid to be used for generating 3D reconstructions efficiently and seamlessly with just one model. The joint appearance and geometry prior captured in Orchid can be used for various downstream tasks. Notably, Orchid enables joint generation of images, depth, and surface normal from an given text prompt using a single model. Sampling from Orchid with a single color image as condition produces accurate and consistent depth and normals, which rival SOTA monocular depth and normal prediction models trained specifically for those tasks. It also excels in joint inpainting of color, depth, and surface normals when used as an unconditional diffusion prior, a capability that is unique to Orchid. We anticipate Orchid will pave the way for advancements in tasks like new-view synthesis from single and sparse observations, densification from sparse depth, and solving inverse problems that entangle appearance and geometry.
@@
In this appendix, we provide additional details on our datasets, model architecture, ablations and training methodology. We also provide additional qualitative results from Orchid for text conditioned color-depth-normal generation as well as image conditioned depth and inpainting tasks, including comparisons to more baselines. Also see the project website containing 3D reconstructions using predictions from Orchid (more details in Section C.5), and qualitative comparisons.
@@ A.1 Architecture
For our VAE, we use a convolutional encoder and decoder with a latent dimension of 8, with $8\times$ spatial downsampling. The VAE has 7 input channels: 3 for RGB, 1 for depth, and 3 for surface normals.
Once the VAE is trained, we keep it frozen when training the latent diffusion model. The latent diffusion model itself is a UNet transformer similar to Stable Diffusion which is conditioned on both time and text embeddings. It has approximately 2B parameters.
@@ A.2 Training
We use a combination of different RGB, depth, and normal losses when training the VAE, weighted differently as explained in Section 3.1 of our paper. Here we provide the values of the weights we found to work best for our setting. For $L_{\mathbf{x}}$, we use $w^{\mathbf{x}}_{1}=1,w^{\mathbf{x}}_{2}=0.1,w^{\mathbf{x}}_{3}=0.1,w^{\mathbf{x}}_{4}=1$. For $L_{\mathbf{d}}$, we use $w^{\mathbf{d}}_{1}=1,w^{\mathbf{d}}_{2}=0.5$. We use $w^{\mathbf{n}}=1$ for $L_{\mathbf{n}}$.
@@
A.3 Dataset
We provide details of the dataset we used for VAE and LDM training in Table 5.
@@ A.3 Dataset
Appendix B Additional ablations

### v2

@@ 3.3 Color-conditioned depth and normal prediction
Figure 4: Text-conditioned generation: Color, depth and normal images generated jointly by our model from a single text prompt. The input text prompt is provided on the top. The generated color, depth, and normal are presented from top to bottom. Results show that Orchid can produce consistent appearance and geometry using its unified LDM for a wide range of text prompts such as close-up object shots or views from complex layouts of indoor and outdoor environment.
To further use Orchid’s learned color-depth-normal prior to jointly generate depth and surface normals from an input color image condition, we append an input color latent $\mathbf{z}^{\mathbf{x}}$ from the pretrained frozen color-only VAE as a condition signal for diffusion through $f_{\theta_{\text{LDM}}}(\mathbf{z}_{t},t;\mathbf{z}^{\mathbf{x}})$. While this is in line with previous work on using diffusion priors for dense prediction [24, 18], our joint latent space allows direct generation of both depth and surface normals from color conditions. When used as an image-conditioned geometry estimator at inference, we drop text conditions (Figure 3b). The sampling procedure is detailed in Alg. 2.
Algorithm 2 Color-conditioned depth-normal sampling
$\mathbf{z}^{\mathbf{x}}$ (color latent condition), $T$ (denoising steps)
@@ 4.1 Implementation details
We use a convolutional encoder and decoder with a latent dimension of 8 for our joint VAE, initialized from a VAE pretrained on datasets of color images alone. Our LDM is a transformer UNet similar to Stable Diffusion [44], pretrained on a large image-text dataset. The diffusion model uses v-prediction. We train both models on 512x512 image resolution. When training our text and color-conditioned diffusion models, we use a DDPM noise schedule [20] with 1000 steps. For diffusion sampling, we adopt a DDIM schedule with 100 steps for text-conditioned generation and 50 steps for color-conditioned generation. We use the zero terminal SNR schedule from [32].
Dataset: We combine many real and synthetic datasets with ground truth depth and surface normals to train our VAE: Omnidata [9], Virtual KITTI [14], Hypersim [42], and DIODE [53]. This amounts to about 2.5M images. Additionally, we also distill predictions from pretrained teacher depth [64] and surface normal [21] models on a text-image dataset with about 100M images.
@@ 4.2 Text-conditioned generation
Orchid is a first-of-its-kind model that can jointly generate color images along with their depth and normal from text prompts. We provide qualitative results of color, depth and normal generated by Orchid from a diverse set of text prompts in Figure 4. Assuming approximate intrinsics, we can use the predicted color, depth and surface normal to optimize 3D Gaussians using Gaussian Splatting [25], as shown in Figure 1. More results (including a user study) are provided in the Appendix.
@@
Table 2: Zero-shot monocular depth estimation: Comparison of our zero shot affine-invariant depth estimation accuracy to that of other methods. Being a joint depth-normal predictor, ours is comparable to the SOTA depth-only prediction model [64] while being better than the other joint depth-normal method [13]. The first, second and third ranking methods are highlighted.
@@
Table 3: Zero-shot monocular normals estimation: Orchid’s color-conditioned surface normal predictions are significantly better than other SOTA methods, including specialized surface normal prediction models. First, second and third ranking methods are highlighted.
Figure 6: Monocular depth prediction: Orchid achieves clearly better performance than Marigold (MG) [24] on far-range objects (row 1), thin structures (row 3), and depth-ambiguous scenes (row 5). It is comparable to Depth-Anything v2 (DA) [64], while being slightly better on very long range objects (row 2). Note that performance on far-away objects is not reflected in Table 2 due to ground truth range limits.
Figure 7: Monocular surface normal prediction: A comparison of our jointly generated normal to that of Marigold [24] and DSINE [2]. Orchid produces more robust results on curved and deformable surfaces.
Dataset DA + DSINE Marigold [24] GeoWizard [13] Orchid
NYUv2 0.059 0.102 0.122 0.040
KITTI 0.144 0.146 0.324 0.082
Table 4: Depth-normal inconsistency: We show the mean error $e_{\text{depth\_normal}}$ ($\downarrow$) in the table. Orchid’s color-conditioned joint depth and normal estimates are significantly more consistent than those produced by GeoWizard, separate Marigold models, or a combination of Depth-Anything V2 and DSINE.
@@
Figure 9: Joint color-depth-normal inpainting: Given a masked region in color-depth-normal data (left column shows masked color), our model inpaints them jointly, leveraging appearance and geometry cues. The baseline uses 3 diffusion models to first inpaint color, then inpaint depth and normals conditioned on it. Unlike ours, it produces geometric inconsistencies e.g. discontinuous edges.
Model Text-conditioned Image-conditioned
CLIP ($\uparrow$) LPIPS ($\uparrow$) $\delta_{1}$ ($\uparrow$) $11.25\degree$ ($\uparrow$)
Orchid 0.316 0.764 96.9 60.6
w/o joint latents 0.312 0.769 96.0 54.8
w/o distillation loss 0.309 0.752 96.1 54.6
w/o pseudo labels - - 96.6 57.9
Table 5: Ablations: We find that our choices of joint latents, distillation loss, and depth-normal pseudo-labels are crucial.
@@ 4.3 Monocular depth and normal prediction
We finetune the text-conditioned Orchid for color-conditioned prediction as explained in Section 3.3). We evaluate the accuracy of depth and normals jointly estimated by Orchid from color images. We follow the protocol in recent approaches and evaluate its zero-shot performance on depth and normal benchmarks we do not train on. In addition, we present qualitative results on web images in Figure 5. As illustrated in the figure, our model can produce high-fidelity and consistent depth and normals from in-the-wild color images from different domains: single objects, and complex indoor and outdoor outdoor scenes.
@@
We evaluate the accuracy of the affine-invariant depth estimated by Orchid on four held-out datasets: NYUDepthv2 [36], KITTI [15], ETH3D [50] and ScanNet [6], against other zero-shot depth estimation baselines. This includes feed-forward [41, 21, 64] as well as diffusion-based models [24, 13]. All baselines (except [13]) are trained solely for depth prediction while Orchid jointly predicts depth and normals. From Table 2, Orchid’s depth estimates are comparable to the state-of-the-art method [64], and better than other diffusion-based baselines. Notably, Orchid achieves better monocular depth estimation performance than Geowizard [13], the only baseline that jointly predicts depth and surface normals. Figure 6 illustrates that our method produces more accurate depth than Marigold [24] on far-range objects, small objects, and depth-ambiguous scenes. These benchmarks do not reflect performance on far-away objects on outdoor datasets like KITTI due to limited ground truth range, where Orchid’s predictions are qualitatively better.
@@
We evaluate Orchid’s normal estimation accuracy on four held-out datasets: NYUv2 [36], ScanNet [6], iBims-1 [27] and Sintel [4]. From Table 3, Orchid is significantly better than baselines at estimating surface normals. This highlights the significance of training a VAE that explicitly encodes color, depth, and normals on a large dataset, as opposed to other diffusion baselines [24, 13, 18] that re-use a frozen color-space VAE for normals. We qualitatively compare to Marigold-normal [24] and DSINE [2] in Figure 7, where our method has better normal predictions, especially on curved and deformable surfaces.
@@
We evaluate the consistency of the depth and surface normal produced by Orchid and compare it to other baselines. We align affine-invariant depth $\mathbf{\hat{d}}$ with ground truth to obtain metric depth $\mathbf{d}$, and compute a 3D pointcloud $\mathbf{p}$ using camera intrinsics. We estimate normals $\mathbf{\hat{n}}=\nabla_{x}\mathbf{p}\times\nabla_{y}\mathbf{p}$ from the pointcloud, and compute its inconsistency $e_{\text{depth\_normal}}=(1-\mathbf{\hat{n}}\cdot\mathbf{n})/2$ with the estimated normals $\mathbf{n}$. As shown in Table 4, we find that Orchid is able to leverage the joint latent space to predict depth and normal that are significantly more consistent. We visualize the depth-normal inconsistency error map in Figure 8. These results clearly demonstrate the benefit of our joint diffusion model in downstream applications.
@@ 4.4 Joint color-depth-normal inpainting
A unique ability of Orchid’s joint color-depth-normal diffusion prior is sampling from the joint distribution conditioned on partial observations, such as jointly inpainting color-depth-normal images. Given color, depth and normals for a frame, we mask all of them for the region to be inpainted. Following the approach from RePaint [34], we use Orchid as an unconditional inpainting prior to generate inpainted regions, as shown in Figure 9. We find that our model is able to generate consistent and realistic completions, without any additional training. A baseline inspired from existing 3D inpainting work [51, 59, 35] is to first run a Stable-Diffusion [44] inpainting model, followed by depth inpatining [54] and normal prediction [18]. Figure 9 shows that this approach is clearly worse, as the color inpainting model produces geometric inconsistencies that are inherited by the geometry prediction models. Please refer to the Appendix for more results.
@@ 4.5 Ablations
We find some of our model and training design choices to be crucial to Orchid’s performance, as shown in Table 5:
Pseudo depth-normal labels: We use a large text-image dataset with depth and normals derived from pretrained teacher models. We find that this significantly boosts the accuracy of generated depth and normals, as measured on color-conditioned prediction. This is consistent with previous depth models that use teacher predictions [16, 63].
Joint latent space: We train an alternative unified color-depth-normal diffusion model that uses separate disentangled latents through a shared diffusion UNet, with an increased input/output dimension to accomodate all three latents. We find this to be significantly worse in terms of text-adherance as seen by a worse CLIP similarity. This is likely due to the joint distribution of disentangled latents drifting significantly from the LDM’s color-pretrained latent space.
Distillation loss: It is used during VAE training to encourage our VAE’s latents to be close in distribution to the original LDM’s latents. We find that dropping it also causes a degradation in the generated images and geometry.
@@ 5 Conclusion
Our work introduces Orchid, a novel joint appearance and geometry diffusion prior that encodes color, depth, and surface normals in a unified latent space. The joint appearance and geometry prior makes Orchid well-suited to various 3D reconstruction applications. Notably, Orchid enables joint generation of images, depth, and surface normal from a text prompt in a single diffusion process. Sampling from Orchid with a color image condition produces accurate and consistent depth and normals, which rival SOTA monocular depth and normal prediction models trained specifically for those tasks. It also excels in joint inpainting of color, depth, and surface normals when used as an unconditional diffusion prior, a capability that is unique to Orchid. We anticipate Orchid will pave the way for advancements in tasks like new-view synthesis from single and sparse observations, densification from sparse depth, and solving inverse problems that entangle appearance and geometry.
Acknowledgements: We are very grateful to James Hays for his encouragement and feedback. We also thank Kyle Genova, Songyou Peng, Thomas Funkhouser, and Leonidas Guibas for their valuable insights during our discussions.
@@
In this appendix, we provide additional details of our datasets, model architecture, ablations, and training methodology. We also provide a runtime analysis, user-study, and additional qualitative results from Orchid for text conditioned color-depth-normal generation as well as image conditioned depth-normal prediction and joint inpainting tasks, including comparisons to more baselines. We conclude with a discussion of our limitations, and scope for future work. Novel-view synthesis videos of 3D reconstructions using predictions from Orchid are provided on the web page https://orchid3d.github.io, along with a discussion in Section E.5.
@@ A.1 Architecture
For our VAE, we use a convolutional encoder and decoder with a latent dimension of 8, with $8\times$ spatial downsampling. The VAE has 7 input channels: 3 for RGB, 1 for depth, and 3 for surface normals. The discriminator (used only during VAE training) is a small ConvNet + MLP.
Once the VAE is trained, we keep it frozen when training the latent diffusion model. The latent diffusion model itself is a UNet transformer similar to Stable Diffusion [43] which is conditioned on both time and text embeddings. It has approximately 2B parameters.
@@ A.2 Training
We use a combination of RGB, depth, and normal losses when training the VAE, with weights as explained in Section 3.1 of our paper. Here, we provide the values of the weights we used for our model. For $L_{\mathbf{x}}$, we use $w^{\mathbf{x}}_{1}=1,w^{\mathbf{x}}_{2}=0.1,w^{\mathbf{x}}_{3}=0.1,w^{\mathbf{x}}_{4}=1$. For $L_{\mathbf{d}}$, we use $w^{\mathbf{d}}_{1}=1,w^{\mathbf{d}}_{2}=0.5$. We use $w^{\mathbf{n}}=1$ for $L_{\mathbf{n}}$. We also use $w^{distill}=10^{-}6$ for $L_{distill}$ and $w^{KL}=10^{-}3$ for $L_{KL}$. Our choice of loss components and their weights for $L_{\mathbf{x}}$ and $L_{\mathbf{KL}}$ are based on standard training recipes for VAEs used in latent diffusion models. For losses we introduce, i.e, $L_{\mathbf{d}}$, $L_{\mathbf{n}}$, and $L_{distill}$, we obtained similar results with weights of similar orders of magnitude, but dropping them completely worsens quality (as shown in our ablations).
@@
A.3 Dataset construction
@@
We provide details of the dataset we use for VAE and LDM training in Table 6. When training the VAE, we sample more heavily from the high-quality real world datasets, rather than our dataset with teacher model predictions. Whereas for the text-conditional LDM training, we sample more heavily from the distillation dataset which contains text-captions. For image-conditioned LDM finetuning, we ignore the text captions, and sample from both real-world and distillation data. While predictions from teacher models are not perfect, models distilled from multiple teachers have performed better in previous work[55]. We remove a few rare examples where depth and normal teacher models disgree (high depth-normal inconsistency) for significant parts of the image.
Appendix B Runtime analysis
Model Diff + Diff Diff + FF Orchid
Inference time (s / img) 4.2* 1.3* 1.2
Table 7: Runtime analysis: Orchid is the fastest way to generate color, depth, and normals. A fair runtime comparison is hard since these methods vary in memory usage. Baselines using multiple models (*) cannot store all models on a GPU and need added weight I/O time that is not included here.
We provide an analysis of the runtime taken for the different approaches discussed in Table 1 in our main paper in Table 7. We report inference times for on a single H100 for all three methods. Our joint generation of color, depth, and normals is significantly faster than generating them with 3 different diffusion models. It is also faster than using discriminative models for depth and normals after an image diffusion process - 1.2 vs 1.3 s per image. Although this difference may seem less significant, please note that we do not include the time taken to move model parameters to/from the GPU, which is required when using multiple models. This I/O time is significantly greater than the inference time for discriminative models.
Appendix C Ablation details